# Proper maps of annuli

### v1

@@ 2. Basic results on proper maps of annuli
If $f\colon\mathbb{C}^{n}\dashrightarrow\mathbb{C}^{N}$ is a rational map so that $f(S^{2n-1})\subset S^{2N-1}$. Then $k_{f}\leq N-1$.
@@
Writing $P=q^{-1}(0)$ for the pole set of $f$, we therefore have that $f(\Sigma_{w}\setminus P)\subset\Sigma_{f(w)}$ for all $w$ where $f(w)$ is defined. As any affine hyperplane that does not go through the origin can be written as $\{z:\langle z,w\rangle=1\}$ for some $w\in\mathbb{C}^{n}$. Therefore, there is an open set of affine hyperplanes which $f$ takes into hyperplanes. Therefore $k_{f}\leq N-1$. ∎
The $k_{f}$ is also invariant under conjugation by linear fractional transformations. The proof is obvious.
@@
Next we consider the $k_{f}$ integer for juxtapositions. We can think of the set of rational ball (or annulus) maps as a convex set; the set of the squared norms $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}$ where $f$ is a map of balls (or annuli) is in fact convex, and the juxtaposition of $f$ and $g$ has the squared norm $(1-t)\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}+t\mathopen{}\mathclose{{\left\lVert g(z)}}\right\rVert^{2}$ that is, it is a convex combination, see for example [dangelo-2009-complexity]. The next proposition therefore shows that for rational ball (or annulus) maps, $k_{f}=N_{f}-1$ happens only on the non-extremal points of this convex set.
@@
Suppose that $F_{t}$ is a juxtaposition of two rational proper maps of balls (or annuli) $F_{0}$ and $F_{1}$, that is, $F_{t}=\sqrt{1-t}\,F_{0}\oplus\sqrt{t}\,F_{1}$ for some $t\in(0,1)$, and further suppose that $\mathopen{}\mathclose{{\left\lVert F_{0}(z)}}\right\rVert^{2}\not\equiv\mathopen{}\mathclose{{\left\lVert F_{1}(z)}}\right\rVert^{2}$. Then $k_{f}\leq N_{f}-2$.
@@
We note that if $k_{F_{0}}\leq N_{F_{0}}-2$ or $k_{F_{1}}\leq N_{F_{1}}-2$, then we are done trivially. Therefore suppose that $k_{F_{0}}=N_{F_{0}}-1$ and $k_{F_{1}}=N_{F_{1}}-1$. Also note from the proof of Proposition˜2.9 that the hyperplanes that contain $f(H)$ are Segre varieties of the sphere and the normal vectors defining these hyperplanes are $f(w)$ as $w$ varies. In particular, we find these vectors in an $N_{f}$-dimensional hyperplane which we can assume is the entire target space. Apart from these facts that we can simply treat these as rational maps. The integers $k_{f}$ and $N_{f}$ and the fact about the normal vectors are all invariant under any LFT and so we only need to prove the result for $F_{0}\oplus F_{1}$. We can also assume that $F_{0}$ has $N_{F_{0}}$ components and $F_{1}$ has $N_{F_{1}}$ components. We can further homogenize both maps up to the same degree and instead of affine hyperplanes we deal with linear subspaces. The dimensions increase by 1, but that is not relevant to us as we are interested in their difference. In other words, what we have is two homogeneous polynomial maps of the same degree $F_{0}$ and $F_{1}$, we have that the components of $F_{0}$ are linearly independent and the components of $F_{1}$ are linearly independent. We further have that for any codimension 1 subspace $H$, the span of $F_{0}(H)$ is of codimension 1 (in the codomain of $F_{0}$) and same for $F_{1}(H)$. It is possible that components of $F_{0}$ and $F_{1}$ are not linearly independent when put together. After composing with a linear we can write $F_{0}=G\oplus\widetilde{F}_{0}$ and $F_{1}=G\oplus\widetilde{F}_{1}$ so that $P=G\oplus\widetilde{F}_{0}\oplus\widetilde{F}_{1}$ has linearly independent components. Write the target space as $\mathbb{C}^{j}\times\mathbb{C}^{k}\times\mathbb{C}^{\ell}$. The fact that $1\leq N_{F_{0}}-k_{F_{0}}$ implies there is a nonzero constant vector $v=(v_{1},v_{2},0)$ ($v_{1}\in\mathbb{C}^{j}$, $v_{2}\in\mathbb{C}^{k}$, $0\in\mathbb{C}^{\ell}$) so that $\xi\cdot v=0$ for all $\xi\in P(H)$. There is also a vector $w=(w_{1},0,w_{3})$ so that $\xi\cdot w=0$ for all $\xi\in P(H)$. We find that unless $v_{2}$ and $w_{3}$ are zero and $v_{1}$ and $w_{1}$ are multiples of each other, we found that $P(H)$ lives in a codimension 2 space and we are done. Thus suppose that there exists a $v=w=(v_{1},0,0)$ for every subspace $H$. But our assumption about the normal vectors of the hyperplanes would give a contradiction. Thus $v$ and $w$ are always linearly independent and we find that $N_{P}-k_{p}\geq 2$. ∎
We have seen that the homogeneous map $H_{d}$ is an annulus map. We also have that juxtapositions of homogeneous maps and their affine embeddings are annulus maps. Here is an example that is not constructed in this way from homogeneous maps.
Example 2.
With the construction introduced in [helal-2025-proper], one finds the following map
(16) $f(z_{1},z_{2})=\mathopen{}\mathclose{{\left(\tfrac{2}{\sqrt{5}}z_{1}^{3},\tfrac{2\sqrt{2}}{\sqrt{5}}z_{1}^{2}z_{2},\tfrac{2}{\sqrt{5}}z_{1}z_{2}^{2},z_{1}z_{2},z_{2}^{2},\tfrac{1}{\sqrt{5}}z_{1}}}\right).$
This map gives a proper map of $\mathbb{A}_{2,\frac{1}{2}}$ to $\mathbb{A}_{6,\frac{1}{4}}$. It takes the unit sphere to the unit sphere and the half sphere to quarter sphere. It takes no other sphere centered at the origin to a sphere centered at the origin. It is not unitarily equivalent to a juxtaposition of several homogeneous maps, nor is it an affine embedding of such a map. If it were, then it would take all spheres centered at the origin to spheres centered at the origin. We have that $N_{f}=6$ and one can compute that the $k_{f}$ number is $3$, so $k_{f}<N_{f}-1$. To see why it is $3$, note that if we restrict to a 1-dimensional affine space $H$, the map becomes a one variable cubic polynomial and so the affine span can be at most $3$. That it is exactly $3$ can be seen by restricting to a hyperplane making $z_{2}$ constant.
We remark that while Theorem˜1.2 implies that for annuli maps, $k_{f}=N-1$ means $f$ is homogeneous, there must be extremal points with $k_{f}<N-1$ in the convex set describing the rational proper annulus maps as the example above shows that there exist maps that are not constructed out of homogeneous maps via juxtapositions.
The construction from [helal-2025-proper] can be used to obtain cubic examples in higher dimensions and also quadratic rational maps with the same properties (not taking more than two zero-centric spheres to zero-centeric spheres, so not constructed out of homogeneous maps).
@@ 3. First gap
Before we prove Theorem˜1.1, we need a result on Hermitian rank that uses similar statement and proof from [dangelo-2012-pfisters]*Theorem 1.1. We only need the quadratic version of this result, but we state the more general result. Recall that the rank of a real polynomial $r(z,\bar{z})$ is the rank of its matrix of coefficients, or alternatively the smallest $k$ such that $r(z,\bar{z})=\sum_{j=1}^{k}\epsilon_{j}\mathopen{}\mathclose{{\left\lvert P_{j}(z)}}\right\rvert^{2}$.
@@
The proof follows by roughly the same argument as in [dangelo-2012-pfisters] which proved the special case when $s_{1}=s_{2}=1$, but with an arbitrary number of factors. What must be newly proved here is a “diagonal version” of the theorem. We only have this proof for two factors, which is sufficient for our purposes. We conjecture that the result holds for any number of factors. Note that the result is not true as stated if the numbers $s_{j}$ have opposite signs. For example, $(\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{2}+1)(\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{2}-1)=\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{4}-1$ is of rank 2 and not $\binom{n+2}{2}=\binom{3}{2}=3$.
@@
The main new claim that must be proved is the “diagonal version” of this theorem. That is, for a real polynomials of real variables $x_{1},\ldots,x_{n}$, we have
@@
where #\# means the number of nonzero coefficients in the given polynomials when they are expanded. We note that if we prove the claim for positive $s_{1},s_{2}$, then it follows for negative $s_{1},s_{2}$ or vice-versa by replacing $x$ by x-x, factoring out a $(-1)^{d}$ and applying the claim for positive $s_{1},s_{2}$.
@@
Next, let us suppose $n=2$. Again let $R(x)=Q(x)(x_{1}+x_{2}+s_{1})(x_{1}+x_{2}+s_{2})$. Write $R(x)=x_{1}^{j}x_{2}^{k}\widetilde{R}(x)$ (and similarly $Q(x)=x_{1}^{j}x_{2}^{k}\widetilde{Q}(x)$) where $\widetilde{R}(x_{1},0)$ and $\widetilde{R}(0,x_{2})$ are both not identically zero. Applying, the one variable version, we find that both $\widetilde{R}(x_{1},0)$ and $\widetilde{R}(0,x_{2})$ have at least $3$ monomials (which are pure powers of $x_{1}$ or $x_{2}$ respectively). One of these terms can be the constant, and therefore, we have $1+2n$ terms in $\widetilde{R}(x)$ which are all pure powers. Next we will show that there is a term in $\widetilde{R}(x)$ that involves both $x_{1}$ and $x_{2}$. Note that the vector field $L=\frac{\partial}{\partial x_{1}}-\frac{\partial}{\partial x_{2}}$ is tangent to each $(x_{1}+x_{2}+s_{j})$. Therefore, if we apply $L$ $\nu$ number of times, we get
@@
As $\widetilde{Q}$ is a polynomial, $L^{\nu}\widetilde{Q}$ must be identically zero for some $\nu$, so take $\nu$ such that $L^{\nu-1}Q$ is not zero but $L^{\nu}Q\equiv 0$. Then $L^{\nu-1}\widetilde{Q}$ is a solution to the transport equation and so it is a function $x_{1}+x_{2}$, that is $L^{\nu-1}\widetilde{Q}(x)=S(x_{1}+x_{2})$ and by the choice of $\nu$ $S$ is not identically zero. But then
@@
is a polynomial in $(x_{1}+x_{2})$ of degree at least 2 and hence it must have a mixed monomial. Since $L^{\nu-1}\widetilde{R}(x)$ has a mixed monomial, then so does $\widetilde{R}(x)$ as $L$ cannot manufacture mixed monomials out of pure ones. Therefore, $R(x)$ has at least $1+2\cdot 2+1=\binom{2+2}{2}=6$ terms.
Moreover, we guarantee that there are at least 3 terms with $x_{1}^{j}x_{2}^{m}$ for $m\geq k$, and at least 3 terms with $x_{1}^{\ell}x_{2}^{m}$ for $\ell>j$ and $m\geq k$. Note that also there are no terms at all with $x_{1}^{\ell}x_{2}^{m}$ for $\ell<j$. Similarly, there are at least 3 terms with $x_{1}^{m}x_{2}^{k}$ for $m\geq j$, and at least 3 terms with $x_{1}^{m}x_{2}^{\ell}$ for $\ell>k$ and $m\geq j$. And no terms with $x_{1}^{m}x_{2}^{\ell}$ for $\ell<k$.
@@
All the monomials of $r_{\mu}(x_{1},x_{2})(x^{\prime})^{\mu}$ must appear in $R(x)$ by minimality of $\mu$. We thus apply the two-dimensional result to find that there must be in fact 6 such monomials, and the distribution of these $a_{\ell m}$ is as in the “Moreover” remark above. We use this this remark for every pair of variables to count monomials. For every pair there is one monomial that we might be also counting for other pairs, so we count it only once. Then there are two monomials corresponding to each variable that may appear in the count for every pair that includes this variable. Finally, for every pair there is at least one monomial that does not appear in the count for every other pair. In other words, We have at least
@@
Now that we have the claim, the theorem follows in the same way as in the proof of the polynomial version of Theorem 2.1 in [dangelo-2012-pfisters]. That is, there exists some ordering of monomials so that if we write out the matrix of coefficients of the expressions involved, multiplication by $\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$ moves along diagonals. The rank of the matrix of coefficients is the rank of the function. The rank of a matrix can be estimated from below by considering the number of nonzero terms in the superdiagonal that is furthest up and to the right that has nonzero terms. Up to multiplication with a monomial, the terms on this diagonal are polynomials in $\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$, and so by replacing $x_{j}=\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$ we obtain the situation of the claim and the proof follows. See [dangelo-2012-pfisters] for more details. ∎
@@
By Theorem˜2.2, $f$ is a rational map that takes $S^{2n-1}$ to $S^{2N-1}$ and $rS^{2n-1}$ to $RS^{2N-1}$. Write $f=\frac{p}{q}$. Lemma 3.11 from [helal-2025-proper] gives us
(25) $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=\Bigl(b_{0}+b_{1}\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1^{2}\bigr)\Bigr)\cdot\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}+Q_{2}(z,\bar{z})\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1^{2}\bigr)\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-r^{2}\bigr)$
for a real polynomial $Q_{2}(z,\bar{z})$.
If $Q_{2}\equiv 0$, then by the argument in the proof of Theorem 3.12 from [helal-2025-proper], $f$ is a polynomial, that is, $q\equiv 1$. Then $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=b_{0}+b_{1}\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)$ is of bidegree $(1,1)$ (degree 1 in $z$ and degree 1 in $\bar{z}$). Therefore, $\deg f=1$. Proposition˜2.7 implies that $f$ is an affine embedding and that $R\geq r$.
Suppose $Q_{2}\not\equiv 0$, we will get a contradiction. By applying Proposition˜2.4, we reduce to the case when $N=N_{f}$. The dimension $N$ is then the number of components of $f$, which is equal to the number of components of $p$, which we can assume are linearly independent as $N_{f}=N$. Therefore, $N$ is equal to the rank of $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$. Theorem˜3.1 gives us
@@
We therefore obtained $N_{f}=n$ and that the degree is $1$. The conclusion follows from Proposition˜2.5 or Proposition˜2.7 depending on whether $N>n$ or $N=n$. The interval is sharp; when $N=\binom{n+1}{2}$, then besides the affine embeddings we can also have the map $H_{2}$ (depending on $R$), which is clearly not unitarily equivalent to an affine embedding. ∎
@@
4. Homogeneous maps via the generic hyperplane rank

### v2

@@ 3. General hyperplane rank and its connection to the $X$-variety
Suppose $f\colon\mathbb{C}^{n}\dashrightarrow\mathbb{C}^{N}$ is a rational map so that $f(S^{2n-1})\subset S^{2N-1}$. Then $k_{f}\leq N-1$.
@@
Writing $P=q^{-1}(0)$ for the pole set of $f$, we therefore have that $f(\Sigma_{w}\setminus P)\subset\Sigma_{f(w)}$ for all $w$ where $f(w)$ is defined. Any affine hyperplane that does not go through the origin can be written as $\{z:\langle z,w\rangle=1\}$ for some $w\in\mathbb{C}^{n}$. Therefore, there is an open set of affine hyperplanes which $f$ takes into hyperplanes. Therefore $k_{f}\leq N-1$. ∎
The general hyperplane rank $k_{f}$ is also invariant under conjugation by linear fractional transformations. The proof is obvious.
@@
Next we consider the $k_{f}$ integer for juxtapositions. We can think of the set of rational ball (or annulus) maps as a convex set; the set of the squared norms $\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}$ where $f$ is a map of balls (or annuli) is in fact convex, and the juxtaposition of $f$ and $g$ has the squared norm $(1-t)\mathopen{}\mathclose{{\left\lVert f(z)}}\right\rVert^{2}+t\mathopen{}\mathclose{{\left\lVert g(z)}}\right\rVert^{2}$ that is, it is a convex combination, see for example [4]. The next proposition therefore shows that for rational ball (or annulus) maps, $k_{f}=N_{f}-1$ happens only on the non-extremal points of this convex set.
@@
Suppose that $F_{t}$ is a juxtaposition of two rational proper maps of balls (or annuli) $F_{0}$ and $F_{1}$, that is, $F_{t}=\sqrt{1-t}\,F_{0}\oplus\sqrt{t}\,F_{1}$ for some $t\in(0,1)$, and further suppose that $\mathopen{}\mathclose{{\left\lVert F_{0}(z)}}\right\rVert^{2}\not\equiv\mathopen{}\mathclose{{\left\lVert F_{1}(z)}}\right\rVert^{2}$. Then $k_{F_{t}}\leq N_{F_{t}}-2$.
@@
We note that if $k_{F_{0}}\leq N_{F_{0}}-2$ or $k_{F_{1}}\leq N_{F_{1}}-2$, then we are done trivially. Therefore suppose that $k_{F_{0}}=N_{F_{0}}-1$ and $k_{F_{1}}=N_{F_{1}}-1$. Also note from the proof of Proposition˜3.2 that the hyperplanes that contain $f(H)$ are Segre varieties of the sphere and the normal vectors defining these hyperplanes are $f(w)$ as $w$ varies. In particular, we find these vectors in an $N_{f}$-dimensional hyperplane, which we can assume is the entire target space. The integers $k_{f}$ and $N_{f}$ and the fact about the normal vectors are all invariant under any LFT and so we only need to prove the result for $F_{0}\oplus F_{1}$. We can also assume that $F_{0}$ has $N_{F_{0}}$ components and $F_{1}$ has $N_{F_{1}}$ components. We can further homogenize both maps up to the same degree and instead of affine hyperplanes we deal with linear subspaces. The dimensions increase by 1, but that is not relevant to us as we are interested in their difference. In other words, what we have is two homogeneous polynomial maps $F_{0}$ and $F_{1}$ of the same degree, and we have that the components of $F_{0}$ are linearly independent and the components of $F_{1}$ are linearly independent. We further have that for any codimension 1 subspace $H$, the span of $F_{0}(H)$ is of codimension 1 (in the codomain of $F_{0}$) and same for $F_{1}(H)$. It is possible that components of $F_{0}$ and $F_{1}$ are not linearly independent when put together. After composing with a linear map, we can write $F_{0}=G\oplus\widetilde{F}_{0}$ and $F_{1}=G\oplus\widetilde{F}_{1}$ so that $P=G\oplus\widetilde{F}_{0}\oplus\widetilde{F}_{1}$ has linearly independent components. Write the target space as $\mathbb{C}^{j}\times\mathbb{C}^{k}\times\mathbb{C}^{\ell}$. The fact that $1\leq N_{F_{0}}-k_{F_{0}}$ implies there is a nonzero constant vector $v=(v_{1},v_{2},0)$ with $v_{1}\in\mathbb{C}^{j}$, $v_{2}\in\mathbb{C}^{k}$, and $0\in\mathbb{C}^{\ell}$ so that $\xi\cdot v=0$ for all $\xi\in P(H)$. There is also a vector $w=(w_{1},0,w_{3})$ so that $\xi\cdot w=0$ for all $\xi\in P(H)$. We find that unless $v_{2}$ and $w_{3}$ are zero and $v_{1}$ and $w_{1}$ are multiples of each other, $P(H)$ lives in a codimension 2 space and we are done. Thus suppose that there exists a $v=w=(v_{1},0,0)$ for every subspace $H$. But our assumption about the normal vectors of the hyperplanes would give a contradiction. Thus $v$ and $w$ are always linearly independent and we find that $N_{P}-k_{P}\geq 2$. ∎
@@ 4. First gap
Before we prove Theorem˜1.1, we need a result on Hermitian rank that uses similar statement and proof from [5]*Theorem 1.1. We only need the quadratic version of this result, but we state the more general result. Recall that the rank of a real polynomial $r(z,\bar{z})$ is the rank of its matrix of coefficients, or alternatively the smallest $k$ such that $r(z,\bar{z})=\sum_{j=1}^{k}\epsilon_{j}\mathopen{}\mathclose{{\left\lvert P_{j}(z)}}\right\rvert^{2}$.
@@
The proof follows by roughly the same argument as in [5] which proved the special case when $s_{1}=s_{2}=1$, but with an arbitrary number of factors. What must be newly proved here is a “diagonal version” of the theorem. We only have this proof for two factors, which is sufficient for our purposes. We conjecture that the result holds for any number of factors. Note that the result is not true as stated if the numbers $s_{j}$ have opposite signs. For example, $(\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{2}+1)(\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{2}-1)=\mathopen{}\mathclose{{\left\lvert z}}\right\rvert^{4}-1$ is of rank 2 and not $\binom{n+2}{2}=\binom{3}{2}=3$.
@@
The main new claim that must be proved is the “diagonal version” of this theorem. That is, for real polynomials of real variables $x_{1},\ldots,x_{n}$, we have
@@
where #\# means the number of nonzero coefficients in the given polynomials when they are expanded. We note that if we prove the claim for positive $s_{1},s_{2}$, then it follows for negative $s_{1},s_{2}$ or vice-versa by replacing $x$ by x-x, factoring out a $(-1)^{2}$ and applying the claim for positive $s_{1},s_{2}$.
@@
Next, let us suppose $n=2$. Again let $R(x)=Q(x)(x_{1}+x_{2}+s_{1})(x_{1}+x_{2}+s_{2})$. Write $R(x)=x_{1}^{j}x_{2}^{k}\widetilde{R}(x)$ (and similarly $Q(x)=x_{1}^{j}x_{2}^{k}\widetilde{Q}(x)$) where $\widetilde{R}(x_{1},0)$ and $\widetilde{R}(0,x_{2})$ are both not identically zero. Applying the one variable version, we find that both $\widetilde{R}(x_{1},0)$ and $\widetilde{R}(0,x_{2})$ have at least $3$ monomials (which are pure powers of $x_{1}$ or $x_{2}$ respectively). One of these terms can be the constant, and therefore, we have $1+2n$ terms in $\widetilde{R}(x)$ which are all pure powers. Next we will show that there is a term in $\widetilde{R}(x)$ that involves both $x_{1}$ and $x_{2}$. Note that the vector field $L=\frac{\partial}{\partial x_{1}}-\frac{\partial}{\partial x_{2}}$ is tangent to each $(x_{1}+x_{2}+s_{j})$. Therefore, if we apply $L$ $\nu$ number of times, we get
@@
As $\widetilde{Q}$ is a polynomial, $L^{\nu}\widetilde{Q}$ must be identically zero for some $\nu$, so take $\nu$ such that $L^{\nu-1}Q$ is not zero but $L^{\nu}Q\equiv 0$. Then $L^{\nu-1}\widetilde{Q}$ is a solution to the transport equation $Lu=0$, and so it is a function of $x_{1}+x_{2}$, that is, $L^{\nu-1}\widetilde{Q}(x)=S(x_{1}+x_{2})$ and by the choice of $\nu$, $S$ is not identically zero. But then
@@
is a polynomial in $(x_{1}+x_{2})$ of degree at least 2 and hence it must have a mixed monomial. Since $L^{\nu-1}\widetilde{R}(x)$ has a mixed monomial, then so does $\widetilde{R}(x)$ as $L$ cannot manufacture mixed monomials out of pure ones. Therefore, $R(x)$ has at least $1+2\cdot 2+1=6=\binom{2+2}{2}$ terms. In particular, we have the following terms:
Statement 1.
There are at least 3 terms with $x_{1}^{j}x_{2}^{m}$ for $m\geq k$. There are at least 3 terms with $x_{1}^{m}x_{2}^{k}$ for $m\geq j$. There is at least one term with $x_{1}^{\ell}x_{2}^{m}$ for $\ell>j$ and $m>k$. There are no terms with $x_{1}^{\ell}x_{2}^{m}$ for $\ell<j$ or $m<k$.
@@
All the monomials of $r_{\mu}(x_{1},x_{2})(x^{\prime})^{\mu}$ must appear in $R(x)$ by minimality of $\mu$. We thus apply the two-dimensional result to find that there must be in fact 6 such monomials, and the distribution of these monomials is given in ˜1. We use this result for every pair of variables to count monomials. For every pair there is one monomial that we might be also counting for other pairs, so we count it only once. Then there are two monomials corresponding to each variable that may appear in the count for every pair that includes this variable. Finally, for every pair there is at least one monomial that does not appear in the count for every other pair. In other words, we have at least
@@
Now that we have the claim, the theorem follows in the same way as in the proof of the polynomial version of Theorem 2.1 in [5]. That is, there exists some ordering of monomials so that if we write out the matrix of coefficients of the expressions involved, multiplication by $\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$ moves along diagonals. The rank of the matrix of coefficients is the rank of the function. The rank of a matrix can be estimated from below by considering the number of nonzero terms in the superdiagonal that is furthest up and to the right that has nonzero terms. Up to multiplication with a monomial, the terms on this diagonal are polynomials in $\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$, and so by replacing $x_{j}=\mathopen{}\mathclose{{\left\lvert z_{j}}}\right\rvert^{2}$ we obtain the situation of the claim and the proof follows. See [5] for more details. ∎
@@
By Theorem˜2.2, $f$ is a rational map that takes $S^{2n-1}$ to $S^{2N-1}$ and $rS^{2n-1}$ to $RS^{2N-1}$. Write $f=\frac{p}{q}$ in lowest terms. By Lemma˜2.5 we have
(24) $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=\Bigl(1+b\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)\Bigr)\mathopen{}\mathclose{{\left\lvert q(z)}}\right\rvert^{2}+Q(z,\bar{z})\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-r^{2}\bigr)$
for a real polynomial $Q(z,\bar{z})$ and $b=\frac{1-R^{2}}{1-r^{2}}$.
If $Q\equiv 0$, then since $\frac{p}{q}$ is in lowest terms, we must have $q$ is constant, so we may assume $q\equiv 1$. Then $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}=1+b\bigl(\mathopen{}\mathclose{{\left\lVert z}}\right\rVert^{2}-1\bigr)$ is of bidegree $(1,1)$ (degree 1 in $z$ and degree 1 in $\bar{z}$). Therefore, $\deg f=1$. Proposition˜2.8 implies that $f$ is an affine embedding and that $R\geq r$.
Suppose $Q\not\equiv 0$, we will get a contradiction. By applying Proposition˜2.4, we reduce to the case when $N=N_{f}$. The dimension $N$ is then the number of components of $f$, which is equal to the number of components of $p$, which we can assume are linearly independent as $N_{f}=N$. Therefore, $N$ is equal to the rank of $\mathopen{}\mathclose{{\left\lVert p(z)}}\right\rVert^{2}$. Theorem˜4.1 gives us
@@
We therefore obtained $N_{f}=n$ and that the degree is $1$. The conclusion follows from Proposition˜2.6 or Proposition˜2.8 depending on whether $N>n$ or $N=n$. The interval is sharp; when $N=\binom{n+1}{2}$, besides the affine embeddings we can also have the map $H_{2}$ (depending on $R$), which is clearly not unitarily equivalent to an affine embedding. ∎
@@
5. Homogeneous maps via the general hyperplane rank